# Collision-Free Humanoid Traversal in Cluttered Indoor Scenes

### v1

@@ I Introduction
While legged locomotion in complex environments has seen remarkable advances for quadrupeds [9, 19, 38, 4, 22, 18, 28, 35, 11, 23, 41, 27, 2] and humanoids [20, 42, 10, 30, 21, 25, 13, 34, 31, 1, 40], existing works are often limited in their ability to handle traversal in cluttered indoor scenes (full-spatial obstacle layouts and intricate, realistic geometries), as shown in Table I. These limitations collectively point to the lack of an effective representation for humanoid–obstacle relationships during collision avoidance: (i) existing works [42, 1, 35, 11, 23, 41, 27, 2] typically obtain penalty signals only when collisions occur, yielding sparse and delayed supervision. This forces reinforcement learning (RL) to depend on inefficient trial-and-error exploration, thus calling for a representation that can provide anticipatory and dense guidance; (ii) conventional representations expose the policy to raw, high-dimensional environmental measurement independently of humanoid–obstacle spatial relationships, forcing the policy to infer traversal decisions through implicit kinematic reasoning.
@@ II Related Works
Figure 1: Overall pipeline. We learn a visuomotor policy that maps diverse obstacle geometries and spatial layouts to corresponding whole-body traversal skills. Left: HumanoidPF for whole-body traversal learning. (Top) Construction of HumanoidPF, a reformulation of APF tailored for humanoid whole-body traversal; (Bottom) its use as informative perceptual representation and collision-avoidance rewards. Right: Scalable training and deployment pipeline. (Top) Hybrid scene generation for constructing diverse and challenging training environments; (Middle) parallel training of multiple specialist policies followed by distillation into a single generalist policy; (Bottom) sim-to-real deployment via Click-and-Traverse, enabling intuitive loco-navigation teleoperation in cluttered indoor scenes. Sections III-A,  III-B and  III-C provide detailed descriptions of the HumanoidPF for traversal learning, the scalable training, and deployment pipeline, respectively.
@@ III-B1 Hybrid scene generation
We observe that highly challenging obstacle layouts constitute merely a long-tail subset in most existing datasets [5, deitke2022️, 24], since typical indoor scenes feature orderly object arrangements and clearly delineated walkable regions. Simply scaling up the dataset does not alleviate this problem. Therefore, we propose a novel hybrid scene generation scheme that enriches realistic 3D indoor datasets with procedurally synthesized ”extreme” obstacles, where full-spatial constraints are jointly present.

### v2

@@ I Introduction
While legged locomotion in complex environments has seen remarkable advances for quadrupeds [12, 23, 8, 44, 25, 22, 43, 33, 40, 15, 28, 47, 32, 3, 14, 5, 27] and humanoids [24, 13, 35, 6, 48, 26, 30, 17, 39, 36, 1, 2, 46], existing works are often limited in their ability to handle traversal in cluttered indoor scenes (full-spatial obstacle layouts and intricate, realistic geometries), as shown in Table I. These limitations collectively point to the lack of an effective representation for humanoid–obstacle relationships during collision avoidance: (i) existing works [5, 27, 15, 14, 48, 2, 40, 28, 47, 32, 3] typically obtain penalty signals only when collisions occur, yielding sparse and delayed supervision. This forces reinforcement learning (RL) to depend on inefficient trial-and-error exploration, thus calling for a representation that can provide anticipatory and dense guidance; (ii) conventional representations expose the policy to raw, high-dimensional environmental measurement independently of humanoid–obstacle spatial relationships, forcing the policy to infer traversal decisions through implicit kinematic reasoning.
@@ II Related Works
Figure 1: Overall pipeline. We learn a visuomotor policy that maps diverse obstacle geometries and spatial layouts to corresponding whole-body traversal skills. Left: HumanoidPF for whole-body traversal learning. (Top) Construction of HumanoidPF, a reformulation of APF tailored for humanoid whole-body traversal; (Bottom) its use as informative perceptual representation and collision-avoidance rewards. Right: Scalable training and deployment pipeline. (Top) Hybrid scene generation for constructing diverse and challenging training environments; (Middle) parallel training of multiple specialist policies followed by distillation into a single generalist policy; (Bottom) sim-to-real deployment via “Click-and-Traverse”, an intuitive loco-navigation teleoperation in cluttered indoor scenes. Sections III-A,  III-B and  III-C provide detailed descriptions of the HumanoidPF for traversal learning, the scalable training, and deployment pipeline, respectively.
@@ III-B1 Hybrid scene generation
We observe that highly challenging obstacle layouts constitute merely a long-tail subset in most existing datasets [7, deitke2022️, 29], since typical indoor scenes feature orderly object arrangements and clearly delineated walkable regions. Simply scaling up the dataset does not alleviate this problem. Therefore, we propose a novel hybrid scene generation scheme that enriches realistic 3D indoor datasets with procedurally synthesized ”extreme” obstacles, where full-spatial constraints are jointly present.